OpenStudy (anonymous):

Which is the equation of a line that is parallel to 2x+5y=4 and goes through the point (5,-4) A) Y= -2/5x-6 B)Y= -2/5x+2 C)Y= 2/5x-6 D)Y= -2/5x-2

OpenStudy (danjs):

ok, it would probably be easiest to put the equation into the slope intercept form to begin y = mx + b, where m is the slope of the line

OpenStudy (anonymous):

5y=2x+4?

OpenStudy (danjs):

almost, you want to subtract 2x on both sides, so you will have a -2x

OpenStudy (danjs):

5y = -2x + 4

OpenStudy (anonymous):

oh yeah lol okay got it. Now what?

OpenStudy (danjs):

divide everything by 5

OpenStudy (danjs):

you want y = .....

OpenStudy (anonymous):

y=-2/5-2?

OpenStudy (danjs):

where did the x go? and 4 divided by 5 is, 4/5

OpenStudy (anonymous):

i accident divded the 4 by -2 lol

OpenStudy (danjs):

thats ok, so this is what you get for the given line in the slope intercept form \[y = \frac{ -2 }{ 5 }x + \frac{ 4 }{ 5 }\]

OpenStudy (anonymous):

But my pretest doesn't give me that for any of my answers.. :(

OpenStudy (danjs):

that is just the given line, we now need to use information from that to find a different line with the point they gave you

OpenStudy (anonymous):

oh okie dokie how does one do that?

OpenStudy (danjs):

The key point here, is that the slope of a parallel line , will be the same slope m = -2/5 in the y = m*x + b for y = (-2/5)x + 4/5

OpenStudy (danjs):

we now have the slope (-2/5) and a point (x,y) = (5,-4) given in the problem.

OpenStudy (danjs):

Point-Slope form of a line, Given slope: m = -2/5 Point : (x1, y1) = (5, -4) Fill in for this point slope equation: y - y1 = m*(x - x1)

OpenStudy (anonymous):

I got 9=m(-7) is that riht?

OpenStudy (danjs):

Point - slope form of a line y - y1 = m*(x - x1) put in: m = -2/5 x1 = 5 y1 = -4

OpenStudy (danjs):

y - (-4) = (-2/5)(x - 5)

OpenStudy (danjs):

all i did was put the numbers in the equation

OpenStudy (danjs):

You can now simplify that, and it should look like one of the answers, a line parallel to the given line

OpenStudy (danjs):

y - (-4) is the same as y + 4 y + 4 = (-2/5)*(x-5)

OpenStudy (danjs):

distributing the -2/5 into the (x-5) you get \[y + 4 = \frac{ -2 }{ 5 }x + 5\frac{ 2 }{ 5 }\]

OpenStudy (danjs):

\[y + 4 = \frac{ -2 }{ 5 }x + 2\]

OpenStudy (danjs):

take away 4 from both sides \[y = \frac{ -2 }{ 5 }x -2\]

OpenStudy (anonymous):

so y=-2/5-6?

OpenStudy (anonymous):

no -2

OpenStudy (anonymous):

-2/5-2

OpenStudy (danjs):

(-2/5)x - 2

OpenStudy (anonymous):

yass! thank you bro!

OpenStudy (danjs):

The things you need to remember: Slope intercept: Y = m*X+b m= slope, b = y-axis intercept Point -slope form y - y1 = m(x-x1) m=slope , (x1,y1) is a point on the line A line parallel to a given line will have the same slope, m